# Chenxu Wang

### v5

@@ 2.1 Training and Testing
In terms of data handling, all datasets were standardized before being utilized for model training and testing. This preprocessing included removing duplicates, imputing missing values, and selecting relevant features. Specifically, SMILES and InChIKeys were standardized using appropriate tools to ensure consistency in the representation of compound structures. Additionally, the log S values were normalized prior to model input to mitigate the impact of numerical variations on model training.
@@ 2.2 Data Preprocessing
In our study, we employed the RDKit library to convert SMILES (Simplified Molecular Input Line Entry System) strings into molecular objects. These molecular objects serve as the foundational data structure for further extraction and analysis of molecular features. RDKit is a widely used cheminformatics toolkit that allows for the processing and manipulation of chemical information, and it was essential for our feature extraction pipeline.
@@ 2.4 Molecular Feature Extraction
Feature extraction is a pivotal component in our model. We used one-hot encoding to convert all categorical variables into binary zeros and ones. Elemental symbols were encoded into 66-bit one-hot vectors, and hybridization features were encoded using 7-bit one-hot vectors. This approach not only streamlines the handling of categorical variables by machine learning algorithms but also maintains feature independence and prevents model biases due to changes in feature ordering[29].
Utilizing the RDKit toolkit, we meticulously extracted diverse features of atoms and chemical bonds from SMILES representations of molecules and tabulated these features (refer to Tables 1 and 2). On this foundation, atoms and chemical bonds were conceptualized as nodes and edges in graph theory, respectively. These graph features were then formatted for deep learning applications using the PyG (PyTorch Geometric) library, with node features having a dimensionality of 92 and edge features having a dimensionality of 10. This comprehensive feature extraction and encoding procedure not only precisely maps the molecular structure but also equips the deep learning model with a well-suited composite description of atoms and chemical bonds.
@@ 2.6 Graph Convolution Network
In our research, we introduced a single-layer Graph Convolutional Network (GCN) architecture to handle graph data. This GCN utilizes convolution operations on the graph structure to adeptly capture the relationships between the features of nodes and their adjacent neighbors. In our model, the GCN layer principally serves to transform node features from 92 dimensions to 128 dimensions, thus augmenting their representational power. Through this approach, the GCN layer not only enhances our comprehension of the nodes’ local structures but also provides a higher-dimensional feature representation to facilitate the analysis of complex interactions within the graph data, strongly supporting subsequent analytical and predictive endeavors[30].
@@ 2.9 YZS-Model
The initial step of the model involves utilizing a Graph Convolutional Network (GCN) to process the graphical representation of drug molecules, effectively revealing the spatial structural features of the molecules. By analyzing the interactions between the atoms and their chemical bonds, the GCN layer can abstract complex feature vectors that describe the molecular topology, laying a solid foundation for in-depth structural analysis. After converting SMILES to a graph representation, where the nodes represent the feature vectors of each atom and the edges represent the bonds, the model uses a GCN to process this graphical representation of drug molecules. This effectively reveals the spatial structural features of the molecules. By analyzing the interactions between the atoms and their chemical bonds, the GCN layer abstracts complex feature vectors that describe the molecular topology, providing a robust basis for detailed structural analysis.
The model iteratively extracts features for each atom within the molecule over k iterations, obtaining the features within each molecule’s k-neighborhood. These final molecular structure feature sequences are used as inputs to the Transformer, providing a solid foundation for subsequent solubility predictions. This process effectively captures complex interactions and long-range dependencies within molecules, thereby enhancing the model’s predictive accuracy.
@@
Subsequently, the model employs a Transformer architecture to further refine the feature output from the GCN layer. Utilizing self-attention mechanisms, the Transformer can discern the broad dependencies among atoms within the molecule, even if they are spatially distant in the molecular structure. This stage of analysis not only focuses on the microstructural details of the molecule but also captures the comprehensive internal interactions, forming a sophisticated feature representation that encompasses extensive global information. Atoms within a molecule can interact even if they are far apart, which can further influence solubility. Therefore, the Transformer architecture is used to analyze molecular features from a global perspective. By leveraging self-attention mechanisms, the Transformer can identify broad dependencies among atoms within the molecule, even if they are spatially distant in the molecular structure. This stage of analysis not only focuses on the microstructural details but also captures comprehensive internal interactions, forming a sophisticated feature representation that encompasses extensive global information.
Furthermore, the LSTM layer is used to deepen the understanding of the features generated by the Transformer architecture. Given the potential temporal dependencies inherent in molecular structure sequences, the LSTM effectively captures the temporal information in the sequence data through its unique gating mechanisms. This processing stage is crucial for revealing the dynamic nature of molecular features across sequences, providing deep temporal insights for solubility prediction.
@@
Finally, through a fully connected layer, the model transforms the output from the LSTM layer into the final solubility prediction. These layers not only synthesize high-dimensional features from the GCN, Transformer, and LSTM but also link these features to the solubility of drug molecules through nonlinear transformations. The entire model is designed to effectively map the spatial and temporal characteristics of molecules to their physicochemical property of solubility, ensuring high accuracy and reliability in the solubility predictions for organic drugs.
@@ 3.2 Interpretative Analysis of YZS-Model
Molecular Complexity and Connectivity: Formal Charge: The presence of formal charges generally increases molecular polarity, thus enhancing solubility. Degree of Connectivity: Higher connectivity indicates more complex molecular structures. For instance, more hydrogen atoms can increase hydrophobicity, affecting solubility[38].
Structural Properties: Aromaticity: Aromatic rings with polar functional groups (e.g., hydroxyl, carboxyl, or amino groups) can enhance solubility through hydrogen bonding with water. Hybridization (sp2 vs. sp3): Sp3 hybridization tends to form more complex, non-planar structures compared to sp2, increasing intermolecular voids and potential polar interactions, thus improving solubility[39].

### v6

@@ 2.1 Training and Testing
In terms of data processing, all datasets were standardized prior to model training and testing. This included steps such as removing duplicates, shuffling the molecules, and feature selection. Specifically, SMILES and InChIKeys were matched one-to-one. We crawled InChIKeys information from the ACS database and compared it with SMILES to prevent mismatches. In the ablation experiments, due to the involvement of training set ablation, to prevent the concentration of any specific feature, we reshuffled the data based on a normal distribution. Additionally, the log S data, as collected by the original compilers, had already undergone normalization to avoid numerical discrepancies that could impact model training.
@@ 2.2 Data Preprocessing
In our study, we employed the RDKit library to convert SMILES (Simplified Molecular Input Line Entry System) strings into molecular objects.Influenced by Ahmad’s paper[21],these molecular objects serve as the foundational data structure for further extraction and analysis of molecular features. RDKit is a widely used cheminformatics toolkit that allows for the processing and manipulation of chemical information, and it was essential for our feature extraction pipeline.
@@ 2.4 Molecular Feature Extraction
In this study’s model, the feature extraction step is crucial. Influenced by Ahmad’s paper[21] on feature extraction, we utilized the RDKit tool to convert the molecular SMILES representation into Mol objects. We then traversed all atoms and chemical bonds, using RDKit’s member methods to accurately extract various features of atoms and bonds. These features were organized into tables (see Table 1 and Table 2). On this basis, atoms and bonds were abstracted as nodes and edges in graph theory, and these graph features were converted into a format suitable for deep learning using the PyG library, with edges and vertices represented by their respective high-dimensional feature vectors. The node feature vectors have a dimension of 92, while the edge feature vectors have a dimension of 10.
We extracted the majority of features present in organic compounds to serve as model features. Since these feature representations cannot be directly processed by the model, we employed one-hot encoding to encode all multi-type variables, converting them into binary forms of zero or one. Specifically, element symbols are encoded using 66-bit one-hot vectors, which cover the vast majority of atoms that constitute organic compounds. Additionally, hybridization features are encoded using 7-bit one-hot vectors, encompassing the most common hybridization forms found in organic materials. This encoding strategy not only simplifies the processing of categorical variables by machine learning algorithms but also ensures the independence of features and effectively prevents model bias caused by changes in feature order [29], thereby streamlining the model’s feature extraction and utilization process.
@@ 2.6 Graph Convolution Network
Based on a comparison of various existing GNN model architectures, we introduced one of the classic applications of GNN, the Graph Convolutional Network (GCN) architecture, to process graph data. GCN utilizes convolution operations on graph structures to gradually and effectively capture the feature relationships between nodes and their neighbors. In this study, by applying the GCN architecture, we can fully capture the molecular features within the k-radius neighborhood of each atom. This approach fully leverages the information within the graph structure of organic molecules, providing robust support for the prediction and analysis of molecular properties.[30].
@@ 2.9 YZS-Model
After converting SMILES into graph representations, the nodes of the graph correspond to the feature vectors of each atom, while the edges represent the bonds. Therefore, the first step of the model involves processing the graphical representation of drug molecules using a Graph Convolutional Network (GCN), which effectively reveals the spatial structural features of the molecules. By analyzing the interactions between atoms and their chemical bonds, the GCN layer abstracts information that describes the molecular topology. It outputs aggregated feature vectors for each atom within the k-radius neighborhood, which are then combined into a feature sequence and passed to the next module. This approach not only captures the overall molecular information but also considers the influence of neighboring atoms.
@@
Within a molecule, atoms can interact with each other even if they are relatively distant, potentially affecting solubility. Therefore, a Transformer architecture is used to analyze molecular features from a global perspective. Based on the feature sequence provided by the previous module, the Transformer leverages the self-attention mechanism to identify extensive dependencies between atoms within the molecule, regardless of their spatial distance in the molecular structure. This stage of analysis focuses on the intricate details of the molecular microstructure and captures the comprehensive internal interactions, resulting in an advanced feature representation that encompasses extensive global information. The module ultimately outputs a tensor sequence containing the attention-distributed molecular structure sequence and features, which is then handed over to the LSTM for further processing.
Subsequently, the LSTM layer is employed to deepen the understanding of the features generated by the Transformer architecture. Specifically, some features produced by the Transformer may have little or no impact on the solubility of organic compounds. The LSTM layer selectively forgets these less impactful features, retaining only the more significant ones, thus enhancing the model’s ability to extract important information. Additionally, considering that the molecular structure sequence output by the Transformer might contain temporal dependencies, and the Transformer’s mechanism tends to be less effective in capturing such dependencies, the LSTM’s unique gated mechanism effectively captures the temporal dependencies within the sequence data. The output sequence from the final LSTM layer is then passed to the next stage. This processing phase is crucial for revealing the dynamic characteristics of molecular features as they change over the sequence, providing deep temporal insights for solubility prediction.
@@
Finally, a fully connected layer translates the output of the LSTM layer into the final solubility prediction value. This layer not only integrates the high-dimensional features from the GCN, Transformer, and LSTM but also projects these features linearly to relate them to the solubility of drug molecules. The design of the entire model aims to achieve an effective mapping from the spatial and temporal characteristics of the molecules to their physicochemical property of solubility, ensuring high accuracy and reliability in predicting the solubility of organic drugs.
@@ 3.2 Interpretative Analysis of YZS-Model
Molecular Complexity and Connectivity: Formal Charge: Formal charge refers to the unequal sharing of electrons between atoms in some covalent bonds, resulting in some atoms within the molecule carrying partial positive charges while others bear partial negative charges. This distribution of charge increases the polarity of the molecule, thereby enhancing its solubility[tsiper2001charge]. Degree of Connectivity: Connectivity refers to the number of bonds between atoms in a molecule. Higher connectivity indicates a more complex molecular structure, with more atomic connections. An increase in connectivity generally leads to increased molecular complexity as more atoms are interlinked to form larger molecular structures. More complex molecular structures may form more interactions within the solvent, which can affect the solubility of the molecule in the solvent. For instance, more complex structures may increase hydrophobicity, thus reducing solubility [38].
Structural Properties: Aromaticity: Molecules with aromaticity typically possess a ring-shaped conjugated system, commonly found in aromatic compounds such as benzene rings. Functional groups with significant polarity, such as hydroxyl (-OH), carboxyl (-COOH), or amino (-NH2) groups, can induce polar interactions on an aromatic ring, forming hydrogen bonds with water and enhancing the molecule’s interaction with water, thereby improving solubility[bounds1981polarization]. Hybridization (sp2 vs. sp3): sp3 hybridization tends to form more complex, non-planar structures that may increase intermolecular voids, providing more opportunities for polar interactions and thus improving solubility [39].